The Bradley Corporation produces a product with the following costs as of July 1, 20X1: Material $5 per unit Labor 3 per unit Overhead 1 per unit Beginning inventory at these costs on July 1 was 3,200 units. From July 1 to December 1, 20X1, Bradley produced 12,400 units. These units had a material cost of $4, labor of $6, and overhead of $4 per unit. Bradley uses LIFO inventory accounting. a. Assuming that Bradley sold 13,800 units during the last six months of the year at $19 each, what is its gross profit? b. What is the value of ending inventory?
Answer #1

LIFO method assumes that the inventory purchased/produced later is sold first

Cost per unit of beginning inventory , 3200 units = 5+3+1 = $9 per unit

Cost per unit of units produced during the 6 months, 12400 units = 4+6+4 = $14 per unit

Calculation of Gross Profit:

Sales 13,800*19

262,200

Less: COGS (12,400*14+1,400*9)

186,200

Gross profit

$76,000

Value of ending inventory = 1800 units @ $9 = $16,200